OpenStudy (anonymous):

Which of the following is not a way to represent the solution of the inequality 2(x - 2) <= 2? A number line with a closed circle on 3 and shading to the right. A number line with a closed circle on 3 and shading to the left. 3 >= x x <= 3

OpenStudy (anonymous):

2(x-2) <= 2 2x - 4 <= 2 +4 +4 2x <= 6 2 2 x = 3

OpenStudy (anonymous):

x >= 3

OpenStudy (phi):

This part is ok: 2x <= 6 2 2 x <= 3 (Notice that you keep the <=. You changed it to =, and then on your next post switched it to x>=3, which is wrong) Notice the last 2 answers 3 >= x x <= 3 mean the same thing (3 greater than or equal to x means x is less than or equal to 3) So you have to choose which of the first two answes is WRONG.